which equation represents a line that is paralle to the line represented by the equation y=3/5x+5/3?
Answer by Alan3354(69443)

which equation represents a line that is paralle to the line represented by the equation y=3/5x+5/3?
-----------------
Assuming you mean (3/5)x, or 3x/5 and not 3/(5x),
There are an infinite number of lines parallel to that one.
They all have the same slope, m = 3/5.
-----------
They're y = (3/5)x + b, where b can be any number.